OpenStudy (anonymous):

multiply (2*root6)-(4*root5)(3*root6+6*root5)

OpenStudy (mathteacher1729):

Do you mean \[2\sqrt{6}-4\sqrt{5}(3\sqrt{6}+6\sqrt{5})\] or \[(2\sqrt{6}-4\sqrt{5})(3\sqrt{6}+6\sqrt{5})\]?

OpenStudy (anonymous):

the second one

OpenStudy (mathteacher1729):

How far have you gotten in this problem? What about it do you feel confident about? What about it do you feel ... not so confident about? Can you multiply things like \(\sqrt{8}\sqrt{6}\)? Can you multiply things like \(2\sqrt{8}\cdot 4\sqrt{10}\) ? I can give better help if I know a bit more about what you can do now and what is giving you trouble. :)